# Improving the Accuracy of DC Optimal Power Flow Formulations via Parameter Optimization

### v1

@@ Abstract
DC Optimal Power Flow (DC-OPF) problems optimize the generators’ active power setpoints while satisfying constraints based on the DC power flow linearization. The computational tractability advantages of DC-OPF problems come at the expense of inaccuracies relative to AC Optimal Power Flow (AC-OPF) problems which accurately model the nonlinear steady-state behavior of power grids. This paper proposes an algorithm that significantly improves the accuracy of the generators’ active power setpoints from DC-OPF problems with respect to the corresponding AC-OPF problems over a specified range of operating conditions. Using sensitivity information in a machine learning-inspired methodology, this algorithm tunes coefficient and bias parameters in the DC power flow approximation to improve the accuracy of the resulting DC-OPF solutions. Employing the Truncated Newton Conjugate-Gradient (TNC) method—a Quasi-Newton optimization technique—this parameter tuning occurs during an offline training phase, with the resulting parameters then used in online computations. Numerical results underscore the algorithm’s efficacy with accuracy improvements in squared two-norm and $\infty$-norm losses of up to $90\%$ and $79\%$, respectively, relative to traditional DC-OPF formulations.
Index Terms:
DC optimal power flow (DC-OPF), AC-OPF, machine learning, parameter optimization.
Nomenclature
  • $\mathcal{N}$
    Set of buses
  • $\mathcal{E}$
    Set of branches
  • $\mathbf{j}$
    Imaginary unit $\mathbf{j}=\sqrt{-1}$
  • $\mathbf{s}^{\text{d}}$
    Complex power demand; $\mathbf{s}^{\text{d}}=\mathbf{p}^{\text{d}}+\mathbf{j}\mathbf{q}^{\text{d}}$
  • $Y^{s}$
    Bus shunt admittance
  • $Y$
    Complex branch line admittance
  • $Y^{c}$
    Complex branch shunt admittance
  • $\bar{s}$
    Thermal branch limit
  • $\underline{\text{s}}^{\text{g}},\overline{\text{s}}^{\text{g}}$
    Complex power generation bounds
  • $\underline{\text{V}},\overline{\text{V}}$
    Voltage magnitude bounds
  • $\mathbf{z}$
    Branch impedance; $\mathbf{z}{=}\mathbf{r}+\mathbf{j}\mathbf{x}$
  • $\mathbf{A}$
    Branch-bus incidence matrix
  • $\mathbf{s}^{\text{g}}$
    Complex power generation; $\mathbf{s}^{\text{g}}{=}\mathbf{p}^{\text{g}}+\mathbf{j}\mathbf{q}^{\text{g}}$
  • $\mathbf{S}$
    Complex power flow; $\mathbf{S}={\mathbf{P}}+\mathbf{j}{\mathbf{Q}}$
  • $\mathbf{V}$
    Complex voltage; $\mathbf{V}=|\mathbf{V}|\angle\bm{\theta}$
  • $c$
    Active power generation cost vector
  • $\mathbf{b}$
    DC power flow coefficient parameter vector
  • $\bm{\gamma},\bm{\rho}$
    DC power flow bias parameter vectors
@@ I Introduction
Optimal Power Flow (OPF) is a fundamental tool for power system design and operation. OPF problems optimize system performance (e.g., minimizing generation costs or reducing voltage deviations) while satisfying both equality constraints from a power flow model and inequality constraints from operational limits. In addition to significant economic implications [1], OPF problems are at the heart of many algorithms for addressing uncertainty [2], planning infrastructure investments [3], improving stability margins [4], etc.
@@
The AC power flow equations accurately model the steady-state behavior of power systems by relating the complex power injections and line flows with the voltage phasors. This motivates solution of AC Optimal Power Flow (AC-OPF) problems that incorporate the AC power flow equations. However, AC-OPF problems are computationally challenging, with these problems being non-convex and NP-hard in the worst-case [5, 6, 7]. These challenges are exacerbated for extensions of OPF problems that consider discrete characteristics, such as those related to transmission switching [8, 9] and unit commitment [10], as well as uncertainties due to variable renewable energy generation and flexible loads [2].
Since first being formulated by Carpentier in 1962 [11], OPF algorithms have undergone continual development [12, 13, 14, 15]. Most existing methods seek a local optimum due to the nonconvex nature of OPF problems that results from the AC power flow equations. Machine learning (ML) techniques are emerging as tools to address OPF challenges, offering computational speedups but often lacking interpretability and consistency with physical intuition [16].
Due to these challenges, engineers often employ power flow approximations [17] to obtain more tractable convex OPF formulations, sometimes by optimizing the linearization point to account for supply and demand uncertainty [18]. Using a common linear approximation known as the DC power flow [19] yields the DC Optimal Power Flow (DC-OPF) problem, which is widely used for both short- and long-term planning purposes. However, the DC-OPF problem’s tractability advantages come with the drawback of inaccuracies relative to the AC-OPF problem. DC power flow approximations linearize the AC power flow equations by assuming a flat voltage profile, applying a small-angle approximation of the sine function, and neglecting reactive power flows and network losses, thus introducing approximation errors when these assumptions do not hold. Inaccuracies from these assumptions are particularly pronounced when reactive power flows and voltage variations play a significant role, such as in heavily loaded systems or systems with high penetrations of renewable generators. In such cases, the DC-OPF problem may not adequately capture the complex behavior of the power system, leading to suboptimal or infeasible decisions in operational and planning contexts [8, 20, 21, 22].
@@
These limitations motivate enhancements to the DC-OPF problem that aim to improve accuracy while maintaining tractability. The DC power flow approximation is foundational to the DC-OPF problem. This approximation relates active power flow between buses $i$ and $j$, denoted $p_{ij}$, to the phase angle difference $\theta_{i}-\theta_{j}$ through a proportionality coefficient $b_{ij}$: $p_{ij}=b_{ij}(\theta_{i}-\theta_{j})$. While typically derived from line parameters such as resistance $r_{ij}$ and reactance $x_{ij}$ via various heuristics, tailoring these parameter values to specific system and operating range characteristics can yield significant accuracy improvements. Common heuristics choose $b_{ij}=1/x_{ij}$ or $b_{ij}=-\Im(1/(r_{ij}+jx_{ij}))$, where $\Im(\cdot)$ extracts the imaginary part of a complex argument. When resistance is nonzero ($r_{ij}\neq 0$), these choices for $b_{ij}$ produce slightly different results, impacting the accuracy of the DC power flow approximation. Selecting the optimal $b_{ij}$ for a given application is not straightforward and depends on system characteristics and objectives [19].
Additionally, the introduction of bias parameters allows further flexibility by adjusting power injections and flows to account for shunts, HVDC infeeds, phase shifts, and line losses [19]. Two variants of the DC power flow—cold-start and hot-start—differ in their reliance on prior information when determining the $b_{ij}$ coefficients and bias parameters.
@@
Our previous work in [23] introduced an algorithm that optimizes both the coefficient and bias parameters to improve DC power flow accuracy across a range of operating conditions. This optimization selects the DC power flow parameter values which minimize the mismatch between the line flows from the DC and AC power flow models over a training set of scenarios.
@@
While this previous paper focused on optimizing the coefficients and biases within the DC power flow approximation itself [23], specialization of the parameters to specific applications can yield even larger accuracy advantages. Accordingly, this paper extends the concepts from [23] to optimize parameters for the DC power flow approximation in order to improve accuracy of DC-OPF solutions relative to AC-OPF solutions (as opposed to the focus on DC power flow versus AC power flow in our prior work). Specifically, we seek the DC power flow parameters which best reduce the discrepancies of the generator active power setpoints associated with the DC-OPF solutions relative to those from the AC-OPF solutions across a range of load demands. This introduces new challenges since the loss function considers the optimal operational decisions from the DC-OPF problem, resulting in a bilevel structure.
@@
To address these challenges, this paper introduces a new algorithm for the adaptive selection of coefficients $b_{ij}$ and bias parameters within the context of the DC-OPF problem. Building on the approach from our previous work [23], we focus on enhancing the DC-OPF model by leveraging machine learning-inspired techniques for parameter optimization. The algorithm performs offline fine-tuning of $b_{ij}$ and bias parameters using a large training dataset of AC-OPF solutions under various operating conditions. Leveraging differentiable optimization techniques originally developed for deep learning architectures [24, 25], we compute the sensitivities of the DC-OPF solution with respect to these parameters. This enables us to iteratively adjust the parameter values during an offline computation phase by using a gradient-based optimization method, specifically the Truncated Newton Conjugate-Gradient (TNC) algorithm, to minimize the difference between DC-OPF’s generator active power setpoints and their AC-OPF counterparts across the training dataset. The resulting optimized parameter values are then applied to improve the accuracy of the DC-OPF problems during online computations.
@@
  2. 2.
    Computing gradients for the DC-OPF solution with respect to both coefficient and bias parameters to obtain sensitivities of the generator set points as well as our defined loss function.
@@ II Optimal Power Flow Problem
[…]
@@ II-A The AC-OPF Problem
Model 1 presents the AC-OPF formulation. The objective (1a) minimizes the total generation cost by summing the costs for each generator $i\in\mathcal{N}$, modeled as quadratic functions (with coefficients $c_{2i}$, $c_{1i}$, and $c_{0i}$) of the real part of the complex power generation $s_{i}^{\text{g}}$. Constraint (1b) ensures that all voltage magnitudes are within specified limits. Constraint (1c) bounds the complex power outputs of each generator within its feasible operating range, where inequalities on complex quantities are interpreted as bounds on the real and imaginary parts. Constraint (1d) ensures that the power flows into each line terminal do not exceed the line’s thermal limits.
@@
$\min_{s^{\text{g}}_{i},V_{i}}\quad$ $\sum_{i\in\mathcal{N}}c_{2i}(\Re(s^{\text{g}}_{i}))^{2}+c_{1i}\Re(s^{\text{g}}_{i})+c_{0i}$ (1a)
@@
$s^{\text{g}}_{i}-s^{\text{d}}_{i}-(Y^{s}_{i})^{\star}|V_{i}|^{2}=\!\!\sum_{(i,j)\in\mathcal{E}}\!\!S_{ij}+\!\!\sum_{(k,i)\in\mathcal{E}}\!\!S_{ki}$ (1e)
@@
[…]
The constraint (1e) ensures power balance for the power generated, consumed, and flowing through the network at each bus. Constraints (1f) and (1g) represent the power flow in each line of the network. The voltage angle at the reference bus is set to zero in (1h). Finally, constraint (1i) limits the phase angle difference across each line (often used as a proxy for transient stability). The AC-OPF problem is inherently nonlinear and non-convex due to the quadratic nature of the power flow equations. Using nonlinear optimization techniques such as interior-point methods, AC-OPF algorithms frequently find high-quality solutions [14, 12, 13, 15].

### v3

@@ Abstract
DC Optimal Power Flow (DC-OPF) problems optimize the generators’ active power setpoints while satisfying constraints based on the DC power flow linearization. The computational tractability advantages of DC-OPF problems come at the expense of inaccuracies relative to AC Optimal Power Flow (AC-OPF) problems that accurately model the nonlinear steady-state behavior of power grids. This paper proposes an algorithm that significantly improves the accuracy of the generators’ active power setpoints from DC-OPF problems with respect to the corresponding AC-OPF problems over a specified range of operating conditions. Using sensitivity information in a machine learning-inspired methodology, this algorithm tunes coefficient and bias parameters in the DC power flow approximation to improve the accuracy of the resulting DC-OPF solutions. Employing the Truncated Newton Conjugate-Gradient (TNC) method, a Quasi-Newton optimization technique, this parameter tuning occurs during an offline training phase, with the resulting parameters then used in online computations. Numerical results underscore the algorithm’s efficacy with accuracy improvements in squared two-norm and $\infty$-norm losses of up to $90\%$ and $79\%$, respectively, relative to traditional DC-OPF formulations.
@@ I Introduction
Optimal Power Flow (OPF) is a fundamental tool for power system design and operation. OPF problems optimize system performance (e.g., minimizing generation costs or reducing voltage deviations) while satisfying both equality constraints from a power flow model and inequality constraints from operational limits. Along with economic implications [cain2012history], OPF problems are central to many algorithms for addressing uncertainty [roald2022review], expansion planning [majidi2016stochastic], improving stability margins [abhyankar2017tsopfIandII], etc.
@@
The AC power flow equations accurately model the steady-state behavior of power systems by relating the complex power injections with the voltage phasors. This motivates the solution of AC Optimal Power Flow (AC-OPF) problems that incorporate the AC power flow equations. However, AC-OPF problems are computationally challenging, with these problems being non-convex and NP-hard in the worst case [bienstock2019strong, molzahn2017, bukhsh2013]. These challenges are exacerbated for extensions of OPF problems that consider discrete characteristics, such as those related to transmission switching [barrows2014, vanAckooij2018] (where inaccuracies in DC model parameters can significantly impact the quality of switching decisions [taheri2024ac]) and unit commitment [padhy2004], as well as uncertainties due to variable renewable energy generation and flexible loads [roald2022review].
OPF algorithms have been a subject of research for many decades [Momoh1999-1and2, aravena_molzahn_zhang_et_al-go_intro]. Most existing algorithms seek a local optimum due to the non-convex nature of OPF problems. Machine learning (ML) tools can address OPF challenges, offering computational speedups but often lacking interpretability [^1] In this context, interpretability refers to the ability to directly relate a model’s parameters and variables to the physical principles of the power system, a feature often lacking in more complex, black-box approaches. and consistency with physical intuition [Duchesne2020].
Due to these challenges, engineers often employ power flow approximations to obtain more tractable convex OPF formulations [molzahn2019]. Using a common linear approximation of the AC power flow equations known as the DC power flow [stott2009dc] yields the DC Optimal Power Flow (DC-OPF) problem, which is widely used for both short- and long-term planning purposes. While other advanced linearizations of the AC power flow equations exist [coffrin2016linear, molzahn2019], our focus in this paper is on improving the accuracy of the DC-OPF formulation due to its widespread usage.
The DC power flow is derived under a set of standard assumptions: a flat voltage profile (magnitudes near 1.0 p.u.), small voltage angle differences, and negligible network losses and reactive power flows. These assumptions render the model computationally tractable and are often reasonably accurate for high-voltage transmission systems. However, approximation errors are introduced when these assumptions do not hold. While heuristics exist to mitigate some inaccuracies, such as embedding estimated power losses within the load data, such approaches may not be robust across diverse operating conditions [stott2009dc]. Inaccuracies from the DC power flow assumptions are particularly pronounced when reactive power flows and voltage variations play a significant role, such as in heavily loaded systems [taylor1994power] or systems with high penetrations of renewable generators [milano2013overview]. In such cases, the DC-OPF problem may not adequately capture the complex behavior of the power system, leading to suboptimal or infeasible decisions in operational and planning contexts [barrows2014, dvijotham_molzahn-cdc2016, baker2021solutions, Khodaei2010].
While the standard DC-OPF formulation is widely used [stott2009dc], its accuracy limitations have spurred the development of various alternatives [molzahn2019]. These can be broadly categorized into those that propose new model structures, those that tune parameters of existing models, and those that are purely based on machine learning techniques. A significant body of work focuses on creating more accurate linear approximations of the AC power flow equations, e.g., [Coffrin2014]. The survey in [molzahn2019] catalogs a wide array of these relaxations and approximations. While potentially more accurate, these models typically introduce new mathematical structures that pose significant practical implementation challenges, as the conventional DC-OPF structure is deeply embedded in numerous industry-grade and academic tools for complex tasks like unit commitment and transmission planning. Altering these problems’ fundamental structure would require a significant overhaul of these established tools. Another research direction, more aligned with our philosophy, focuses on enhancing existing models by tuning their coefficients and bias parameters. This includes optimizing the linearization point to account for supply and demand uncertainty [Hohmann2019]. This paradigm of enhancing classical, interpretable models has also proven effective for other approximations, such as the LinDistFlow model for distribution systems [Taheri2024LinDistFlow]. Our prior work used a single-level optimization to tune DC power flow parameters to better match AC power flow results [taheri2023optimizing]. While effective at improving the approximation’s accuracy, it did not directly optimize for the end-goal of improving the OPF solution itself. More recently, black-box machine learning models have been developed to predict AC-OPF solutions with high fidelity [Donti2021], though often at the cost of interpretability.
Our work, in contrast, occupies a middle ground. We leverage techniques from the machine learning domain, specifically differentiable optimization layers [Amos2017, Agrawal2019], but our goal is to enhance a classical, interpretable model. This paper distinguishes itself by intentionally preserving the standard DC-OPF formulation to ensure immediate compatibility with existing tools. Furthermore, we introduce a bilevel optimization framework that represents a significant conceptual advance. Rather than simply matching intermediate power flow values, our method tunes the DC-OPF parameters to directly minimize the error in the final optimal generator setpoints. This solution-focused approach carves a third path between pure machine learning and classical, fixed-parameter models, enhancing accuracy while preserving the interpretability and compatibility of the standard DC-OPF model.
@@
The DC power flow approximation is foundational to the DC-OPF problem. This approximation relates active power flow between buses $i$ and $j$, denoted $p_{ij}$, to the phase angle difference $\theta_{i}-\theta_{j}$ through a proportionality coefficient $b_{ij}$: $p_{ij}=b_{ij}(\theta_{i}-\theta_{j})$. While typically derived from line parameters such as resistance $r_{ij}$ and reactance $x_{ij}$ via various heuristics, tailoring these parameter values to specific system and operating range characteristics can yield significant accuracy improvements. Common heuristics choose $b_{ij}=1/x_{ij}$ or $b_{ij}=-\Im(1/(r_{ij}+\mathrm{j}x_{ij}))$, where $\Im(\cdot)$ extracts the imaginary part of a complex argument. When resistance is nonzero ($r_{ij}\neq 0$), these choices for $b_{ij}$ produce slightly different results, impacting the accuracy of the DC power flow approximation. Selecting the optimal $b_{ij}$ for a given application is not straightforward and depends on system characteristics and objectives [stott2009dc]. Additionally, the introduction of bias parameters allows further flexibility by adjusting power injections and flows to account for shunts, HVDC infeeds, phase shifts, and line losses [stott2009dc]. Two variants of the DC power flow—cold-start and hot-start—differ in their reliance on prior information when determining the $b_{ij}$ coefficients and bias parameters.
@@
Our previous work in [taheri2023optimizing] introduced an algorithm that optimizes both the coefficient and bias parameters to improve DC power flow accuracy across a range of operating conditions. This optimization selects the DC power flow parameter values which minimize the mismatch between the line flows from the DC and AC power flow models over a training set of scenarios. We emphasize that the standard DC power flow parameters (e.g., $b_{ij}=1/x_{ij}$) are themselves derived from simplifying assumptions that may not hold in practice. The goal of this parameter optimization is not to redefine the physical reactance or susceptance of a line. Rather, it is to find an effective set of parameter values that causes the DC power flow approximation to better match the results of the full AC power flow model. As discussed above, the standard DC power flow parameters are themselves derived from simplifying assumptions; we seek a different set of parameters that better serves the purpose of more accurately solving an optimization problem across a range of operating conditions.
@@
While the previous paper [taheri2023optimizing] focused on optimizing the coefficients and biases within the DC power flow approximation itself, specialization of the parameters to specific applications can yield even larger accuracy advantages. Accordingly, this paper optimizes parameters for the DC power flow approximation in order to improve accuracy of DC-OPF solutions relative to AC-OPF solutions (as opposed to the focus on DC power flow versus AC power flow in our prior work). Specifically, we seek the DC power flow parameters that best reduce the discrepancies of the generator active power setpoints associated with the DC-OPF solutions relative to those from the AC-OPF solutions across a range of load demands. This represents a significant conceptual and technical advancement that necessitates reformulating as a bilevel optimization problem that differs substantially from the single-level problem in [taheri2023optimizing]. The upper level seeks to tune the DC-OPF parameters, while the lower level solves the convex DC-OPF problem itself. This enhanced structure introduces new challenges, as the loss function must consider the optimal operational decisions from the lower-level optimization.
@@
Leveraging differentiable optimization techniques developed for deep learning architectures [Amos2017, Agrawal2019], this paper proposes a new algorithm for adaptively selecting coefficients $b_{ij}$ and bias parameters within the specific context of DC-OPF problems. The algorithm performs offline fine-tuning of $b_{ij}$ and bias parameters using a training dataset of AC-OPF solutions under various representative operating conditions. We differentiate the lower-level problem to compute the sensitivities of the DC-OPF solutions with respect to these parameters. This enables us to iteratively adjust the parameter values during an offline phase by using a gradient-based optimization method, specifically the Truncated Newton Conjugate-Gradient (TNC) algorithm, to minimize the difference between DC-OPF’s generator active power setpoints and their AC-OPF counterparts across the training dataset. The resulting optimized parameter values are then applied to improve the accuracy of the DC-OPF problems during online computations.
@@
    Formulating the task of tuning DC-OPF parameters as a bilevel optimization problem, where the goal is to minimize the error in the optimal generator setpoints relative to full AC-OPF solutions.
  2. 2.
@@
    Computing gradients for the DC-OPF solution with respect to both coefficient and bias parameters to obtain sensitivities of the generator setpoints as well as our defined loss function, enabled by modern differentiable programming techniques.
  4. 4.
@@ II Optimal Power Flow Problem
To introduce notation and necessary background material, this section describes the AC-OPF formulation and the DC-OPF linear approximation that simplifies the AC-OPF problem to improve tractability at the cost of accuracy. Buses, lines and generators in the network are denoted by the sets $\mathcal{N}$, $\mathcal{E}$ and $\mathcal{G}$, respectively. Here, $\mathcal{E}$ represents the set of individual branches; we note that multiple parallel branches may exist between two buses, and our formulation treats each one distinctly with its own parameters and flow limits. Each bus $i\in\mathcal{N}$ has a voltage phasor $V_{i}$ with phase angle $\theta_{i}$, along with a shunt admittance $Y_{i}^{S}$, complex power demand $s^{\text{d}}_{i}=\text{p}^{\text{d}}_{i}+\mathrm{j}\text{q}^{\text{d}}_{i}$, and generated complex power $s^{\text{g}}_{i}=p^{\text{g}}_{i}+\mathrm{j}q^{\text{g}}_{i}$, where $\mathrm{j}$ is the imaginary unit $\mathrm{j}=\sqrt{-1}$. Buses without generators are modeled as having upper and lower generation limits of zero. The complex power flows entering each end of line $(j,k)\in\mathcal{E}$ are represented by $S_{jk}$ and $S_{kj}$. Series admittance parameters for each line $(j,k)$ are $Y_{jk}$ and $Y_{kj}$ and the line’s shunt admittance is $Y^{c}_{jk}$. The real and imaginary components of a complex number are indicated by $\Re(\,\cdot\,)$ and $\Im(\,\cdot\,)$, respectively, $\|\cdot\|_{\infty}$ denotes the $L_{\infty}$-norm, and $\|\cdot\|_{2}$ denotes the $L_{2}$-norm. The complex conjugate and the transpose of a matrix are denoted by $(\,\cdot\,)^{\star}$ and $(\,\cdot\,)^{\top}$, respectively, and $\angle(\,\cdot\,)$ is the angle of a complex argument. Upper and lower limits are represented by $(\overline{\,\cdot\,})$ and $(\underline{\,\cdot\,})$.
@@ II-A The AC-OPF Problem
Model 1 presents the AC-OPF formulation. The objective (1a) minimizes the total generation cost by summing the costs for each generator $i\in\mathcal{N}$, modeled as quadratic functions (with coefficients $c_{2i}$, $c_{1i}$, and $c_{0i}$) of the real part of the complex power generation $s_{i}^{\text{g}}$. Following typical modeling practices, we assume that $c_{2i}\geq 0$ such that the generator cost function is convex. Constraint (1b) ensures that all voltage magnitudes are within specified limits. Constraint (1c) bounds the complex power outputs of each generator within its feasible operating range, where inequalities on complex quantities are interpreted as bounds on the real and imaginary parts. Constraint (1d) enforces the line flow limits.
@@
$\operatorname*{minimize}_{s^{\text{g}}_{i},V_{i}}\quad$ $\sum_{i\in\mathcal{N}}c_{2i}(\Re(s^{\text{g}}_{i}))^{2}+c_{1i}\Re(s^{\text{g}}_{i})+c_{0i}$ (1a)
@@
$s^{\text{g}}_{i}-s^{\text{d}}_{i}-(Y^{s}_{i})^{\star}|V_{i}|^{2}=\!\!\sum_{(i,j)\in\mathcal{E}}\!\!S_{ij}+\!\!\sum_{(k,i)\in\mathcal{E}}\!\!S_{ik}$ (1e)
@@
$\textbf{variables}:$ $~s^{\text{g}}_{i}~(\forall i\in\mathcal{N}),V_{i}~(\forall i\in\mathcal{N})$
Constraint (1e) ensures power balance at each bus and (1f) and (1g) represent the power flows on each line. The reference angle is set to zero in (1h). Finally, (1i) limits the phase angle difference across each line. The AC-OPF problem is inherently nonlinear and non-convex due to the quadratic nature of the power flow equations. Using nonlinear optimization techniques such as interior-point methods, AC-OPF algorithms frequently find high-quality solutions [aravena_molzahn_zhang_et_al-go_intro, Momoh1999-1and2, gopinath2022].